### v1

# Solution to a Problem of Erdős Concerning Distances and Points

Abstract.
In 1997, Erdős asked [1] whether for arbitrarily large $n$ there exists a set of $n$ points in $\mathbb{R}^{2}$ that determines $O(\frac{n}{\sqrt{\log n}})$ distinct distances while satisfying the local constraint that every 4-point subset determines at least 3 distinct pairwise distances. We construct $n$-point sets from an $m\times m$ box of the lattice
$L=\{(x,\sqrt{2}y):x,y\in\mathbb{Z}\}\subset\mathbb{R}^{2}.$
The distinct distance bound follows from applying Bernays’ theorem [3] to the number of integers represented by the binary quadratic form $u^{2}+2v^{2}$. The local 4-point constraint is verified through Perucca’s similarity classification [5] of the six similarity types determining exactly two distances.
1. Introduction
Erdős asks whether there exists a set of $n$ points in $\mathbb{R}^{2}$ such that every subset of 44 points determines at least 33 distinct distances, yet the total number of distinct distances is
$\leq C\frac{n}{\sqrt{\log n}}$
for some constant $C>0$ [1], which we write as $O\!\left(\frac{n}{\sqrt{\log n}}\right)$. A formulation also appears as the 659th problem on the Erdős problems website [2].
Our approach uses a standard number-theoretic argument, applying Bernays’ theorem on integers represented by binary quadratic forms to a representation of the lattice’s distinct distances. While the distance counting method is classical, the present note demonstrates that a specific anisotropic variant achieves this global bound while preserving the local constraint. We verify this by a case analysis of the six configurations given by Perucca [5]. Our simple construction proves the following theorem, answering Erdős’ question in the affirmative.
Theorem 1.
For every integer $n\geq 2$ there is a set $P\subset\mathbb{R}^{2}$ with $|P|=n$ satisfying:
  1. (i)
    every 44-point subset of $P$ determines at least 33 distinct pairwise distances; and
  2. (ii)
    $\bigl|\{\|p-q\|:\ p,q\in P,\ p\neq q\}\bigr|=O\!\left(\frac{n}{\sqrt{\log n}}\right)$.
2. The Construction
Consider the anisotropic lattice
$L=\{(x,\sqrt{2}y):x,y\in\mathbb{Z}\}\subset\mathbb{R}^{2}.$
For an integer $m\geq 1$, define the $m\times m$ box subset
$P_{m}=\{(i,\sqrt{2}\,j):0\leq i,j\leq m-1\}.$
Note that $|P_{m}|=m^{2}$.
Definition 2.
For a finite set $P\subset\mathbb{R}^{2}$, let
$D(P):=\{\|p-q\|:p,q\in P,\ p\neq q\}$
denote its set of (nonzero) distinct pairwise distances.
3. Verification of the distinct-distance bound
Since $d\mapsto d^{2}$ is a bijection on the positive reals, we may bound $|D(P_{m})|$ by counting distinct squared distances. For $p=(x_{1},\sqrt{2}\,y_{1})$ and $q=(x_{2},\sqrt{2}\,y_{2})$ in $L$,
$\|p-q\|^{2}=(x_{1}-x_{2})^{2}+2(y_{1}-y_{2})^{2}=u^{2}+2v^{2}$
where $u,v\in\mathbb{Z}$. For $p,q\in P_{m}$ we have $|u|\leq m-1$ and $|v|\leq m-1$, hence
$\|p-q\|^{2}\leq(m-1)^{2}+2(m-1)^{2}<3m^{2}.$
Distinct squared distances in $P_{m}$ are exactly the values of the quadratic form $Q(u,v)=u^{2}+2v^{2}$ with $|u|,|v|\leq m-1$, and as such are bounded by $3m^{2}$. Therefore
$|D(P_{m})|\leq B_{Q}(3m^{2}),$
where $B_{Q}(x)$ denotes the number of positive integers $\leq x$ represented by $Q$.
Theorem 3 (Bernays [3, Eq. (2)]).
Let $f(X,Y)=aX^{2}+bXY+cY^{2}$ be a primitive integral binary quadratic form with non-square discriminant $\Delta=b^{2}-4ac$, and assume $f$ is positive definite. Let $B_{f}(x)$ denote the number of positive integers $\leq x$ represented by $f$. Then there exists a constant $C_{\Delta}>0$ (depending on $\Delta$) such that
$B_{f}(x)\sim C_{\Delta}\,\frac{x}{\sqrt{\log x}}\qquad(x\to\infty).$
The form $Q(u,v)=u^{2}+2v^{2}$ is primitive and positive definite, with discriminant
$\Delta=0^{2}-(4)\cdot(1)\cdot(2)=-8,$
which is not a square. Theorem 3 therefore applies to $Q$, giving
$B_{Q}(x)\sim C_{-8}\,\frac{x}{\sqrt{\log x}}\qquad(x\to\infty).$
In particular,
$|D(P_{m})|\leq B_{Q}(3m^{2})=O\!\left(\frac{m^{2}}{\sqrt{\log m}}\right)=O\!\left(\frac{|P_{m}|}{\sqrt{\log|P_{m}|}}\right).$
Corollary 4.
For every integer $n\geq 2$, there exists a set $P\subset\mathbb{R}^{2}$ with $|P|=n$ such that
$|D(P)|=O\!\left(\frac{n}{\sqrt{\log n}}\right).$
Proof.
Let $m=\lceil\sqrt{n}\rceil$ and choose any $n$-point subset $P\subset P_{m}$. Removing points cannot increase the number of distinct distances, so $|D(P)|\leq|D(P_{m})|$. Also $m^{2}\asymp n$ and $\log m\asymp\log n$, hence
$|D(P)|\leq|D(P_{m})|=O\!\left(\frac{m^{2}}{\sqrt{\log m}}\right)=O\!\left(\frac{n}{\sqrt{\log n}}\right).$
4. Verification of the local 44-point constraint
The local constraint fails if and only if the 4-point set determines exactly two distinct distances. A complete classification of 44-point two-distance configurations (up to similarity) is given in [5] and depicted in Figure 1. Note that the only similarity types that do not contain an equilateral triangle are the square and the regular-pentagon trapezoid.
Figure 1. The six 44-point two-distance configurations (up to similarity). Short vs. long segments are color-coded.
Theorem 5.
For every $m\geq 1$, every 44-point subset of $P_{m}$ determines at least 33 distinct pairwise distances.
Proof.
Let $S\subset P_{m}$ with $|S|=4$. Suppose that $S$ determines exactly two distinct distances. By Perucca’s classification [5], either $S$ is a square, $S$ contains an equilateral triangle, or $S$ is similar to the regular-pentagon trapezoid.
The square case is impossible by Lemma 6, and the equilateral-triangle case is impossible by Lemma 7. The regular-pentagon trapezoid case is impossible by Lemma 8. Hence $S$ cannot determine only two distances and must determine at least three. ∎
4.1. Squares
Lemma 6.
The lattice $L$ contains no non-degenerate square.
Proof.
If four points of $L$ form a square, then the vector between two adjacent vertices is a nonzero vector
$\vec{v}=(u,\sqrt{2}\,v)\in L,\qquad(u,v)\in\mathbb{Z}^{2}\setminus\{(0,0)\},$
and the other side vector is obtained by a $90^{\circ}$ rotation:
$\begin{pmatrix}0&-1\\[2.0pt] 1&0\end{pmatrix}\vec{v}=(-\sqrt{2}\,v,\,u).$
As such, $(-\sqrt{2}\,v,\,u)$ must belong to L, which implies $\sqrt{2}\,v\in\mathbb{Z}$ and $u\in\sqrt{2}\mathbb{Z}$. This is only possible if $\vec{v}=0$ and thus forces a contradiction. ∎
4.2. Equilateral triangles
Lemma 7.
The lattice $L$ contains no non-degenerate equilateral triangle.
Proof.
If $p,q,r\in L$ form an equilateral triangle, then $\vec{v}=q-p\in L$ is nonzero and rpr-p equals $R_{\pm 60}\vec{v}$, where
$R_{60}=\begin{pmatrix}\tfrac{1}{2}&-\tfrac{\sqrt{3}}{2}\\[2.0pt] \tfrac{\sqrt{3}}{2}&\tfrac{1}{2}\end{pmatrix}$.
Write $\vec{v}=(u,\sqrt{2}\,v)$ with $u,v\in\mathbb{Z}$, not both 0. Then
$R_{60}\vec{v}=\left(\frac{u-\sqrt{6}\,v}{2},\ \frac{\sqrt{3}\,u+\sqrt{2}\,v}{2}\right).$
For $R_{60}\vec{v}$ to lie in $L$, there must exist $k\in\mathbb{Z}$ such that
$\frac{\sqrt{3}\,u+\sqrt{2}\,v}{2}=\sqrt{2}\,k,$
so $\sqrt{3}\,u=\sqrt{2}\,(2k-v)\in\mathbb{Q}(\sqrt{2})$. Since $\sqrt{3}\notin\mathbb{Q}(\sqrt{2})$, this forces $u=0$. Then the first coordinate becomes $\frac{\sqrt{6}\,v}{2}$, which is irrational unless $v=0$. This implies $u=v=0$, forcing a contradiction with $\vec{v}\neq 0$. The same argument holds for $R_{-60}$. ∎
4.3. The regular-pentagon trapezoid
Lemma 8.
The lattice $L$ contains no 44-point set similar to the isosceles trapezoid arising from four vertices of a regular pentagon.
Proof.
In this configuration, the ratio of the diagonal length to the side length equals the golden ratio $\varphi=\frac{1+\sqrt{5}}{2}$, hence the ratio of squared lengths (diagonal to side) is $\varphi^{2}=\frac{3+\sqrt{5}}{2}$.
For points in $L$, every squared distance equals $u^{2}+2v^{2}\in\mathbb{Z}$. Therefore any ratio of (nonzero) squared distances determined by a 44-point subset of $L$ is rational, and in particular cannot equal $\varphi^{2}$. ∎
5. Conclusion
We have shown that the anisotropic lattice box $P_{m}\subset\mathbb{Z}\times\sqrt{2}\,\mathbb{Z}$ satisfies both the asymptotic distinct-distance bound
$|D(P_{m})|=O(\frac{|P_{m}|}{\sqrt{\log|P_{m}|}})$
and the local constraint that every 44-point subset determines at least three distinct pairwise distances. For a general $n$, taking any $n$-point subset of $P_{\lceil\sqrt{n}\rceil}$ preserves the local constraint and gives the same asymptotic upper bound (up to constants), as in Corollary 4. The construction ultimately works because $60^{\circ}$ and $90^{\circ}$ rotations, as well as the golden-ratio distance relation, are incompatable with with the quadratic field $\mathbb{Q}(\sqrt{2})$.
Acknowledgements
Gemini 3.0 provided the core ideation behind this approach. The author relied extensively on GPT-5.2 and Gemini 3.0 for drafting, editing, and exploring possible proof strategies. All mathematical claims, proofs, and citations were independently verified by the author, who takes responsibility for any errors. I am grateful to Professor Peter Doyle and Professor Peter Winkler for helpful discussions and feedback on early drafts.